The length of a rectangle is four times the width. If the perimeter of the rectangle is 210 feet find the length.
Answer by Maths68(1474)

Length = l = ?
Width = w
Perimeter = P = 210
Given
Length = 4 * width
l=4w
Perimeter of a Rectangle = length + length + width + width
Perimeter of a Rectangle =2 length +2 width
P = 2l +2 w
Put the values of l,w, p in above equation
210=2(4w)+2w
210=8w+2w
210=10w
Divide by 10 both sides of the equation
210/10=10w/10
21=w
w=21
Width of the rectangle = 21 feet
Length of the rectangle = 4w=4*21=84 feet